OpenStudy (anonymous):

Solve the following system using the substitution method. x + 3y = –5 5x + 6y = –19

OpenStudy (anonymous):

?

jhonyy9 (jhonyy9):

om first you get x=-5-3y if you will put it in the second what you will get ?

jhonyy9 (jhonyy9):

make the calcule

jhonyy9 (jhonyy9):

comen i will help to you

jhonyy9 (jhonyy9):

5(-5-3y)+6y= -19 so can you continue ?

jhonyy9 (jhonyy9):

-25-15y+6y=-19 -9y=6 y=6/-9 y=2/-3 y=-2/3

jhonyy9 (jhonyy9):

x+3(-2/3)=-5 x-6/3=-5 x-2=-5 x=-3

jhonyy9 (jhonyy9):

x=-3 y=-2/3

jhonyy9 (jhonyy9):

this is understandably ?